OpenStudy (anonymous):

multiply:

OpenStudy (anonymous):

\[3x \sqrt{2}*\sqrt{14x}\]

OpenStudy (anonymous):

do you go to connections academy? just wondering

OpenStudy (anonymous):

ummm yes

OpenStudy (anonymous):

ok do you have to solve for x in this problem?

OpenStudy (anonymous):

\[\sqrt{(9x^{2}*2*14x)} = \sqrt{252x^{3}}\]

OpenStudy (anonymous):

I hope this is it you're wishing to get ???

OpenStudy (anonymous):

i think its 6√7x^(3/2)

OpenStudy (anonymous):

@sriramkumar thats not an option for me to choose from

OpenStudy (anonymous):

\[3x \sqrt{2}*\sqrt{14x}\] \[3x \sqrt{28x}\] what two numbers multiplied give you 28 but one having a square root?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

no 4 and 7

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so leave the 7 in the square root and the x as well and take out the 4

OpenStudy (anonymous):

is my answer corrct??

OpenStudy (anonymous):